OpenStudy (anonymous):

simplify .... A= (x^2 + 2x - 24)(x-4)/x-3

jimthompson5910 (jim_thompson5910):

Is the equation \[\large A=\frac{x^2+2x-24}{(x-4)(x-3)}\] ???

OpenStudy (anonymous):

ill give u the whole equation just a sec

OpenStudy (anonymous):

Write an expression to represent the length of the rectangle if the area of the rectangle is x2 + 2x – 24 and the width is (x-4)/ (x-3).

jimthompson5910 (jim_thompson5910):

k, if you can use the equation editor or draw it out

OpenStudy (anonymous):

i cant understand the eqn

jimthompson5910 (jim_thompson5910):

oh I see, So Area = Length * Width This means that A = LW We're given \[\large A=x^2+2x-24\] and \[\large W=\frac{x-4}{x-3}\] So plug these values into A=LW to get \[\large x^2+2x-24=L\times\frac{x-4}{x-3}\] Now multiply both sides by the reciprocal of the width and flip the equation to get \[\large L=(x^2+2x-24)\times\frac{x-3}{x-4}\] Now multiply the fractions to get \[\large L=\frac{(x^2+2x-24)(x-3)}{x-4}\] From there, factor x^2+2x-24 to get (x-4)(x+6), so we now have \[\large L=\frac{ (x-4)(x+6)(x-3)}{x-4}\] which simplifies to \[\large L= (x+6)(x-3)=x^2+3x-18\] So \[\large L=x^2+3x-18\]

OpenStudy (anonymous):

thank u :)